# Towards Open-Vocabulary Audio-Visual Event Localization

### v1

@@ Abstract
The Audio-Visual Event Localization (AVEL) task aims to temporally locate and classify video events that are both audible and visible. Most research in this field assumes a closed-set setting, which restricts these models’ ability to handle test data containing event categories absent (unseen) during training. Recently, a few studies have explored AVEL in an open-set setting, enabling the recognition of unseen events as “unknown”, but without providing category-specific semantics. In this paper, we advance the field by introducing the Open-Vocabulary Audio-Visual Event Localization (OV-AVEL) problem, which requires localizing audio-visual events and predicting explicit categories for both seen and unseen test data at inference. To address this new task, we propose the OV-AVEBench dataset, comprising 24,800 videos across 67 real-life audio-visual scenes (seen:unseen = 46:21), each with manual segment-level annotation. We also establish three evaluation metrics for this task. Moreover, we investigate two baseline approaches, one training-free and one using a further fine-tuning paradigm. Specifically, we utilize the unified multimodal space from the pretrained ImageBind model to extract audio, visual, and textual (event classes) features. The training-free baseline then determines predictions by comparing the consistency of audio-text and visual-text feature similarities. The fine-tuning baseline incorporates lightweight temporal layers to encode temporal relations within the audio and visual modalities, using OV-AVEBench training data for model fine-tuning. We evaluate these baselines on the proposed OV-AVEBench dataset and discuss potential directions for future work in this new field. The dataset and codes will be available at https://github.com/jasongief/OV-AVEL.
@@ 1 Introduction
Audio-visual learning, an essential sub-field of multimodal learning, has garnered increasing attention in recent years, resulting in the development of various research tasks, such as speech recognition [18, 1, 17, 10] and separation [20, 39, 8, 35], sound-source localization [37, 7, 31, 26] and segmentation [55, 27, 33, 5, 15, 59], audio-visual video parsing [42, 12, 19, 56, 57, 58] and question answering [52, 21, 49, 40, 22, 23]. In this paper, we focus on a fundamental research task of Audio-Visual Event Localization (AVEL) [41]. As shown in Fig. 1(a), given a video containing both audio and visual streams, the AVEL task aims to temporally localize segments that contain an audio-visual event (i.e., both audible and visible) and identify its category. For segments that do not satisfy this condition (i.e., only audible/visible or neither), their category is assigned to a special background class. In other words, this task requires perceiving the semantic alignment between audio and visual modalities, known as audio-visual correspondence [2, 3].
@@
Since the pioneering work [41], numerous significant research works have been proposed. For example, these methods aim to improve audio-visual fusion [41, 44, 48, 24, 11, 16, 53], better distinguish the background [47, 54], and localize more precise temporal boundaries [50, 45, 25].
@@ 4.2 Main Results
We propose both a training-free baseline and a fine-tuning baseline to address the OV-AVEL task. In Table 2, we report the performances of both baselines. The training-free baseline achieves 59.2% in Acc., 46.7% in Seg., and 34.0% in Eve., resulting in an average performance of 46.6%. The fine-tuning baseline significantly outperforms the training-free version, showing an 11.2% improvement in the average metric. Moreover, we observe that: the training-free baseline model performs slightly better on the unseen test data; after fine-tuning on training data, the baseline model is improved in recognizing both seen and unseen test data (17.4% and 8.8% in Avg., respectively). The improvement is more pronounced for seen test data because their event classes have been seen in training. However, fine-tuning remains beneficial for unseen test data. We speculate that further fine-tuning helps the model learn temporal relations from training data, facilitating the adaption and updating of prior knowledge from ImageBind to downstream OV-AVEL task. This enables more precise localization of temporal boundaries for unseen test data. This is supported by the event-level metric, which significantly improves from 34.0% to 47.5%. We provide additional evidence and discussions on this in Sec. 4.4. In short, the comparison between the two baselines highlights the benefits of further fine-tuning, especially when some training data with annotations are available. More quantitative (Secs. B, C, D) and qualitative (Sec. F) comparison results are provided in our Appendix .
| Data type | Training-free | Fine-tuning | ||||||
| Acc. | Seg. | Eve. | Avg. | Acc. | Seg. | Eve. | Avg. | |
| total | 59.2 | 46.7 | 34.0 | 46.6 | 67.1 | 56.9 | 49.5 | 57.8 |
| seen | 57.5 | 45.0 | 34.0 | 45.5 | 72.5 | 61.8 | 54.5 | 62.9 |
| unseen | 59.8 | 47.3 | 34.0 | 47.0 | 64.9 | 55.0 | 47.5 | 55.8 |
| Acc. | Seg. | Eve. | Avg. | |
| 1 | 67.1 | 56.9 | 49.5 | 57.8 |
| 2 | 65.4 | 56.0 | 49.2 | 56.9 |
| 3 | 62.8 | 54.0 | 47.3 | 54.7 |
@@ 4.3 Ablation Studies
In this section, we provide some ablation studies on the key configurations adopted in our fine-tuning baseline.
The number of temporal layers . We employ learnable temporal layers to enhance temporal interactions within audio and visual modalities. The results, as shown in Table 3, illustrate the impacts of varying the number of layers. The model achieves the highest average performance using only one temporal layer. Increasing the number of temporal layers may make the model more complex and lead to overfitting, thus degrading the performance. Consequently, we identify to implement the temporal layer, which is lightweight and only introduces 8.4M trainable parameters.
The special class text other. We utilize a special text other to assist the model in classifying events that do not belong to either the seen or unseen classes. We conduct an ablation study to explore its impact. As shown in Table 10, the model using additional other class outperforms that baseline trained without other by 10.8% in average performance. The improvement is consistent across both seen and unseen test data. This underscores the superiority of introducing the additional class text other, which helps prevent the model from misclassifying unknown events or backgrounds as existing seen or unseen classes. In Sec. A of our Appendix, we further show that the employment of other is slightly better than other choices like background.
@@
The strategy for generating . In our fine-tuning baseline, we generate the audio-visual event probability by computing the square root of predicted audio event probability and visual event probability (Eq. 2). We refer to this strategy as Sqrt. Furthermore, we explore two additional variants to obtain . (1) Prob-avg uses the average result of and to generate , i.e., . (2) Fea-avg first generates the fused feature by averaging the audio feature and visual feature , and then computes the cosine similarity (sim) between the fused feature and the text feature , formulated as . We re-train the fine-tuning baseline model using these strategies and evaluate the model on the test set. As shown in Table 5, the Sqrt strategy significantly outperforms the Prob-avg and Fea-avg variants. The geometric mean used by Sqrt is more effective than arithmetic mean at preventing audiovisual event predictions from being misled by a high-probability prediction in one modality. These results indicate the importance of design in generating final audio-visual event probabilities, a factor that future research should also consider.
4.4 Further Discussions
In this section, we provide further results and analyses about our proposed baselines and the studied OV-AVEL task. We hope the findings and discussions inspire future work.
@@ 4.4 Further Discussions
| Cases | Acc. | Seg. | Eve. | Avg. |
| intra only | 67.1 | 56.9 | 49.5 | 57.8 |
| cross only | 54.4 | 45.9 | 39.2 | 46.5 |
| intra + cross | 63.5 | 54.3 | 47.1 | 55.0 |
@@
Intra-modal vs. Cross-modal temporal layers. The temporal layers in our fine-tuning baseline facilitate temporal interactions within the audio and visual modalities (intra-modal). We also attempted to insert some temporal layers to capture cross-modal temporal relations. As shown in Table 7, adding cross-modal temporal layers does not yield improvements. We speculate that the audio and visual features extracted by the pretrained ImageBind model can provide explicit and precise semantics of audio events and visual events, reducing the need for cross-modal interactions. By focusing on the temporal interactions in intra-modality, the model can achieve satisfactory performance.
@@
| Data type | Video-LLaMA2 [9] | Training-free (ours) | ||||||
| Acc. | Seg. | Eve. | Avg. | Acc. | Seg. | Eve. | Avg. | |
| total | 48.9 | 39.1 | 29.8 | 39.3 | 59.2 | 46.7 | 34.0 | 46.6 |
| seen | 50.1 | 40.6 | 32.0 | 40.9 | 57.5 | 45.0 | 34.0 | 45.5 |
| unseen | 48.5 | 38.5 | 29.0 | 38.6 | 59.8 | 47.3 | 34.0 | 47.0 |
Different ratios of training data used for model fine-tuning. As shown in Table 8, we fine-tune the baseline model with various ratios of training data (sampling data for each training class accordingly). Interestingly, we find that the model achieves similar average performance across different data ratios. For instance, using only 25% of the training data, the model performance can reach 57.7% in Avg., close to that achieved with 100% training data. The cost of using less training data is that more training/fine-tuning epochs are required. Additionally, we find that while using 100% of training data is more beneficial for unseen test data recognition, using 50% of training data is the most effective for seen test data recognition. We conjecture that more training data may cause the problem of overfitting for seen classes. Therefore, determining a more balanced training strategy to optimize both seen and unseen data recognition would be a valuable direction for future work.
Comparison with Large Language Model (LLM). Currently, many LLMs are designed for video understanding, though few can handle and analyze both audio and visual inputs simultaneously. We evaluate a recently proposed, open-source audio-visual LLM, Video-LLaMA2 [9], on our OV-AVEBench dataset. Table 9 suggests that the zero-shot performance of Video-LLaMA2 is inferior to our training-free baseline. While Video-LLaMA2 can describe events within entire audio sequence, it faces challenges in achieving segment-level perception of audio-visual alignment. By emphasizing this limitation, we believe that potential LLM-based solutions can be developed for the OV-AVEL task.
@@ 5 Conclusion
Appendix A Further Ablation Study on other
@@
In Sec. 4.3 of our main paper, we have demonstrated that our baseline models using additional class text other outperform models that do not use other. Here, we further compare the employment of other with another option, namely background. The experimental results are shown in Table 10. For both the training-free and fine-tuning baselines, the use of other is slightly better than background. Compared to background, we speculate that the text other can further help the model deal with situations that include other meaningful event classes not listed in the seen and unseen class texts.
Appendix B Comparison between the Training-free Baseline with Another Variant
The training-free baseline introduced in our main paper utilizes ImageBind [14] to extract audio, visual, and textual embeddings. It computes the audio-text and visual-text feature (cosine) similarities to determine final audio-visual event predictions. We refer to this strategy as joint since multimodal features are extracted from a shared feature space. Furthermore, we compare this approach with another variant, where the audio-text and visual-text feature similarities are calculated using feature embeddings from separate backbones. Specifically, for each segment, the pretrained CLAP [46] model is used to extract the audio and text features to generate the audio-text feature similarity; the pretrained CLIP [38] model is used to extract the visual and text features to generate the visual-text feature similarity. Notably, the text encoders of CLAP and CLIP models are different, so the text features are extracted independently. After obtaining the audio-text and visual-text feature similarities, we identify the event categories of the audio segments and visual segments based on the highest similarity values. The final audio-visual event prediction can be made by comparing the consistency of the predicted audio and visual event categories. The experimental results are shown in Table 11. The joint baseline model using ImageBind significantly outperforms the separate variant, with improvements of 4.9%, 3.9%, and 5.2% in the Avg. metric on the total, seen, and unseen test data, respectively. These results indicate the advantages of adopting a joint feature space for multimodal feature embedding, which can better capture semantic alignment among multiple modalities for the OV-AVEL task.
Appendix C Zero-shot Evaluation on AVE [41] Dataset
The AVE dataset is constructed for the closed-set AVEL task [41]. Here, we directly apply our two baseline models to the test set of AVE dataset in a zero-shot inference manner. As shown in Table 12, the fine-tuning baseline continues to outperform the training-free version, demonstrating results competitive with the prior unsupervised state-of-the-art (SOTA) method CMLCL [4]. Notably, CMLCL still uses unlabeled videos of the training set in the AVE dataset. Moreover, if further fine-tuning our baseline model on the AVE dataset, the model can reach 79.6% accuracy without sophisticated designs, approaching the performance of fully-supervised AVEL methods [41, 50, 54, 47, 13]. Nevertheless, we encourage readers to focus on the intrinsic differences: our method is designed for the open-vocabulary AVEL, while prior SOTA methods are tailored specifically for closed-set AVEL.
@@ Appendix C Zero-shot Evaluation on AVE [41] Dataset
Appendix D Class-wise Performance of the Proposed Two Baselines

### v2

@@ Abstract
The Audio-Visual Event Localization (AVEL) task aims to temporally locate and classify video events that are both audible and visible. Most research in this field assumes a closed-set setting, which restricts these models’ ability to handle test data containing event categories absent (unseen) during training. Recently, a few studies have explored AVEL in an open-set setting, enabling the recognition of unseen events as “unknown”, but without providing category-specific semantics. In this paper, we advance the field by introducing the Open-Vocabulary Audio-Visual Event Localization (OV-AVEL) problem, which requires localizing audio-visual events and predicting explicit categories for both seen and unseen test data at inference. To address this new task, we propose the OV-AVEBench dataset, comprising 24,800 videos across 67 real-life audio-visual scenes (seen:unseen = 46:21), each with manual segment-level annotation. We also establish three evaluation metrics for this task. Moreover, we investigate two baseline approaches, one training-free and one using a further fine-tuning paradigm. Specifically, we utilize the unified multimodal space from the pretrained ImageBind model to extract audio, visual, and textual (event classes) features. The training-free baseline then determines predictions by comparing the consistency of audio-text and visual-text feature similarities. The fine-tuning baseline incorporates lightweight temporal layers to encode temporal relations within the audio and visual modalities, using OV-AVEBench training data for model fine-tuning. We evaluate these baselines on the proposed OV-AVEBench dataset and discuss potential directions for future work in this new field.
@@ 1 Introduction
Audio-visual learning, an essential sub-field of multimodal learning, has garnered increasing attention in recent years, resulting in the development of various research tasks, such as sound-source localization [36, 5, 15, 13, 32, 25] and segmentation [53, 55, 16, 26, 27, 12, 14], audio-visual video parsing [40, 48, 9, 54, 56, 57, 50, 6] and generation [38, 28, 18], question answering [19, 47, 20, 21, 22], etc. In this paper, we focus on a fundamental research task of Audio-Visual Event Localization (AVEL) [39]. As shown in Fig. 1(a), given a video containing both audio and visual streams, the AVEL task aims to temporally localize segments that contain an audio-visual event (i.e., both audible and visible) and identify its category. For segments that do not satisfy this condition (i.e., only audible/visible or neither), their category is assigned to a special background class. In other words, this task requires perceiving the semantic alignment between audio and visual modalities, known as audio-visual correspondence [1, 2].
@@
Since the pioneering work [39], numerous significant research works have been proposed. For example, these methods aim to improve audio-visual fusion [39, 42, 46, 23, 8, 58, 17, 51], better distinguish the background [45, 52], and localize more precise temporal boundaries [48, 43, 24].
@@ 4.1 Implementation Details
| Method | seen | unseen | total | |||||||||
| Acc. | Seg. | Eve. | Avg. | Acc. | Seg. | Eve. | Avg. | Acc. | Seg. | Eve. | Avg. | |
| Video-LLaMA2 [7] | 50.1 | 40.6 | 32.0 | 40.9 | 48.5 | 38.5 | 29.0 | 38.6 | 48.9 | 39.1 | 29.8 | 39.3 |
| CLIP [37]&CLAP [44] | 51.4 | 41.4 | 31.9 | 41.6 | 51.6 | 42.2 | 31.6 | 41.8 | 51.5 | 41.9 | 31.7 | 41.7 |
| Training-free (our) | 57.5 | 45.0 | 34.0 | 45.5 | 59.8 | 47.3 | 34.0 | 47.0 | 59.2 | 46.7 | 34.0 | 46.6 |
| CMRA [46] | 65.2 | 58.8 | 54.3 | 59.4 | 36.0 | 31.0 | 26.3 | 31.1 | 44.3 | 38.9 | 34.3 | 39.2 |
| AVE [39] | 76.6 | 63.6 | 56.0 | 65.4 | 44.6 | 33.2 | 24.0 | 34.0 | 53.8 | 41.9 | 33.2 | 42.9 |
| PSP [51] | 75.4 | 66.8 | 61.0 | 67.7 | 33.7 | 28.2 | 24.2 | 28.7 | 45.6 | 39.3 | 34.7 | 39.9 |
| MM-Pyramid [48] | 76.5 | 66.9 | 62.3 | 68.6 | 36.8 | 29.0 | 23.8 | 29.9 | 48.4 | 40.2 | 35.2 | 41.2 |
| Fine-tuning (our) | 72.5 | 61.8 | 54.5 | 62.9 | 64.9 | 55.0 | 47.5 | 55.8 | 67.1 | 56.9 | 49.5 | 57.8 |
@@ 4.2 Main Results
We propose both a training-free (zero-shot) baseline and a fine-tuning baseline to address the new OV-AVEL task. Additionally, we implement several zero-shot and fine-tuning approaches to establish a comprehensive benchmark.
Comparison of Training-free approaches. We compare our training-free baseline with two methods: 1) Video-LLaMA2 [7]. We design task-specific prompts (see our supp. for details) to enable this advanced audio-visual LLM to analyze audiovisual inputs and generate event-prediction texts. 2) CLIP&CLAP. Instead of using ImageBind [11], we employ separate CLIP [37] and CLAP [44] models to extract audio-text and visual-text features, obtaining predictions following Eq. 1. As shown in the upper part of Table 2, our training-free baseline outperforms both methods. While Video-LLaMA2 can describe events within an entire audio sequence, it faces challenges in achieving segment-level perception of audio-visual alignment. Additionally, the comparison with CLIP&CLAP highlights the advantages of using a joint feature space for multimodal feature embedding, which better captures semantic alignment among multiple modalities for the OV-AVEL task.
Comparison of Fine-tuning approaches. We replace the temporal layers in our fine-tuning baseline with core audio-visual fusion modules from prior closed-set AVEL methods (e.g., CMRA [46], AVE [39], PSP [51], MM-Pyr [48]) to enable them perform both seen and unseen event localization for comparison. As shown in the lower part of Table 2, we find that: 1) While complex audio-visual interactions from prior methods may improve seen-class performance, they significantly degrade unseen-class performance, causing their overall results on the total test set to lag far behind ours. 2) Comparing AVE with PSP/MM-Pyr, more advanced interaction modules may exacerbate the imbalance between seen and unseen class recognition, highlighting the challenges of the OV-AVEL task.
Training-free vs. Fine-tuning baselines. Our fine-tuning baseline significantly outperforms the training-free version, showing an 11.2% improvement in the average metric (‘Avg.’) on the total test set. Moreover, we observe that: the training-free baseline model performs slightly better on the unseen test data; after fine-tuning on training data, the baseline model is improved in recognizing both seen and unseen test data (17.4% and 8.8% in Avg., respectively). The improvement is more pronounced for seen test data because their event classes have been seen in training. However, fine-tuning remains beneficial for unseen test data. We speculate that further fine-tuning helps the model learn temporal relations from training data, facilitating the adaption and updating of prior knowledge from ImageBind to downstream OV-AVEL task. This enables more precise localization of temporal boundaries for unseen test data. This is supported by the event-level metric, which significantly improves from 34.0% to 47.5%. We provide additional evidence and discussions on this in Sec. 4.3. In short, the comparison between the two baselines highlights the benefits of further fine-tuning, especially when some training data with annotations are available. More quantitative and qualitative comparison results are provided in our supp. material.
@@ 4.3 Ablation Studies
In this section, we provide some ablation studies on the key configurations adopted in our fine-tuning baseline. More ablations are presented in our supplementary material.
The special class text other. We utilize a special text other to assist the model in classifying events that do not belong to either the seen or unseen classes. We conduct an ablation study to explore its impact. As shown in Table A8, the model using additional other class outperforms that baseline trained without other by 10.8% in average performance. The improvement is consistent across both seen and unseen test data. This underscores the superiority of introducing the additional class text other, which helps prevent the model from misclassifying unknown events or backgrounds as existing seen or unseen classes. In our supp. material, we further show that the employment of other is slightly better than other choices like background.
@@
The strategy for generating . In our fine-tuning baseline, we generate the audio-visual event probability by computing the square root of the product of predicted audio event probability and visual event probability (Eq. 2). We refer to this strategy as Sqrt. Furthermore, we explore two additional variants to obtain . (1) Prob-avg uses the average result of and to generate , i.e., . (2) Fea-avg first generates the fused feature by averaging the audio feature and visual feature , and then computes the cosine similarity (sim) between the fused feature and the text feature , formulated as . We re-train the fine-tuning baseline model using these strategies and evaluate the model on the test set. As shown in Table 4, the Sqrt strategy significantly outperforms the Prob-avg and Fea-avg variants. The geometric mean used by Sqrt is more effective than arithmetic mean at preventing audiovisual event predictions from being misled by a high-probability prediction in one modality. These results indicate the importance of design in generating final audio-visual event probabilities, a factor that future research should also consider.
@@
Different ratios of training data used for model fine-tuning. As shown in Table 6, we fine-tune the baseline model with various ratios of training data (sampling data for each training class accordingly). Interestingly, we find that the model achieves similar average performance across different data ratios. For instance, using only 25% of the training data, the model performance can reach 57.7% in Avg., close to that achieved with 100% training data. Additionally, we find that both 100% and 25% of training data better improve unseen test data recognition, while using 50% of training data is the most effective for seen test data recognition. These results reveal a non-linear link between training data size and the model’s performance, showing more data is not vital for seen/unseen class localization. However, the trade-off of using less training data is the need for more training or fine-tuning epochs. These findings suggest that determining a more balanced training strategy to optimize both seen and unseen data recognition would be a valuable direction for future work.
@@ 5 Conclusion
Acknowledgement This work was supported in part by the National Key R&D Program of China (NO.2024YFB3311602), the National Natural Science Foundation of China (72188101, 62272144, 62020106007, and U20A20183), and the Major Project of Anhui Province (202203a05020011, 2408085J040).
Appendix A The number of temporal layers
Our fine-tuning baseline employs learnable temporal layers to enhance temporal interactions within audio and visual modalities. The results, as shown in Table A7, illustrate the impacts of varying the number of layers. The model achieves the highest average performance using only one temporal layer. Increasing the number of temporal layers may make the model more complex and lead to overfitting, thus degrading the performance. Consequently, we identify to implement the temporal layer, which is lightweight and only introduces 8.4M trainable parameters.
| TF | Data type | w. other | w. background | ||||||
| Acc. | Seg. | Eve. | Avg. | Acc. | Seg. | Eve. | Avg. | ||
| total | 59.2 | 46.7 | 34.0 | 46.6 | 59.1 | 46.6 | 33.8 | 46.5 | |
| seen | 57.5 | 45.0 | 34.0 | 45.5 | 57.5 | 45.1 | 34.0 | 45.5 | |
| unseen | 59.8 | 47.3 | 34.0 | 47.0 | 59.7 | 47.2 | 33.7 | 46.9 | |
| FT | Data type | w. other | w. background | ||||||
| Acc. | Seg. | Eve. | Avg. | Acc. | Seg. | Eve. | Avg. | ||
| total | 67.1 | 56.9 | 49.5 | 57.8 | 66.2 | 56.1 | 48.5 | 56.9 | |
| seen | 72.5 | 61.8 | 54.5 | 62.9 | 71.8 | 60.9 | 53.7 | 62.1 | |
| unseen | 64.9 | 55.0 | 47.5 | 55.8 | 63.9 | 54.1 | 46.4 | 54.8 | |
Appendix B Further Ablation Study on other
@@
In Sec. 4.3 of our main paper, we have demonstrated that our baseline models using additional class text other outperform models that do not use other. Here, we further compare the employment of other with another option, namely background. The experimental results are shown in Table A8. For both the training-free and fine-tuning baselines, the use of other is slightly better than background. Compared to background, we speculate that the text other can further help the model deal with situations that include other meaningful event classes not listed in the seen and unseen class texts.
| Cases | Acc. | Seg. | Eve. | Avg. |
| intra only | 67.1 | 56.9 | 49.5 | 57.8 |
| cross only | 54.4 | 45.9 | 39.2 | 46.5 |
| intra + cross | 63.5 | 54.3 | 47.1 | 55.0 |
Appendix C Intra-modal vs. Cross-modal temporal layers
The temporal layers in our fine-tuning baseline facilitate temporal interactions within the audio and visual modalities (intra-modal). We also attempted to insert some temporal layers to capture cross-modal temporal relations. As shown in Table A9, adding cross-modal temporal layers does not yield improvements. We speculate that the audio and visual features extracted by the pretrained ImageBind model can provide explicit and precise semantics of audio events and visual events, reducing the need for cross-modal interactions. By focusing on the temporal interactions in intra-modality, the model can achieve satisfactory performance.
@@
Appendix D Comparison with the CLIP&CLAP
In sec. 4.2, we compare our training-free baseline with another zero-shot approach, CLIP&CLAP. Here, we provide more implementation details. The training-free baseline introduced in our main paper utilizes ImageBind [11] to extract audio, visual, and textual embeddings. It computes the audio-text and visual-text feature (cosine) similarities to determine final audio-visual event predictions. We refer to this strategy as joint since multimodal features are extracted from a shared feature space. Furthermore, we compare this approach with another variant, where the audio-text and visual-text feature similarities are calculated using feature embeddings from separate backbones. Specifically, for each segment, the pretrained CLAP [44] model is used to extract the audio and text features to generate the audio-text feature similarity; the pretrained CLIP [37] model is used to extract the visual and text features to generate the visual-text feature similarity. Notably, the text encoders of CLAP and CLIP models are different, so the text features are extracted independently. After obtaining the audio-text and visual-text feature similarities, we identify the event categories of the audio segments and visual segments based on the highest similarity values. The final audio-visual event prediction can be made by comparing the consistency of the predicted audio and visual event categories. The experimental results are shown in Table A10. The joint baseline model using ImageBind significantly outperforms the separate variant, with improvements of 4.9%, 3.9%, and 5.2% in the Avg. metric on the total, seen, and unseen test data, respectively. These results indicate the advantages of adopting a joint feature space for multimodal feature embedding, which can better capture semantic alignment among multiple modalities for the OV-AVEL task.
Appendix E Zero-shot Evaluation on AVE [39] Dataset
The AVE dataset is constructed for the closed-set AVEL task [39]. Here, we directly apply our two baseline models to the test set of AVE dataset in a zero-shot inference manner. As shown in Table A11, the fine-tuning baseline continues to outperform the training-free version, demonstrating results competitive with the prior unsupervised state-of-the-art (SOTA) method CMLCL [3]. Notably, CMLCL still uses unlabeled videos of the training set in the AVE dataset. Moreover, if further fine-tuning our baseline model on the AVE dataset, the model can reach 79.6% accuracy without sophisticated designs, approaching the performance of fully-supervised AVEL methods [39, 48, 52, 45, 10]. Nevertheless, we encourage readers to focus on the intrinsic differences: our method is designed for the open-vocabulary AVEL, while prior SOTA methods are tailored specifically for closed-set AVEL.
Appendix F Class-wise Performance of the Proposed Two Baselines